OpenStudy (cenaida):

In the equation 3x + 4y = 9, find x if 2y = 6. A. -2 B. -1 C. 0 D. 1

OpenStudy (freckles):

if 2y=6 then 4y=2(6)=?

OpenStudy (freckles):

then insert that number in place of 4y and solve for 3x then x

OpenStudy (cenaida):

is it 0?

OpenStudy (freckles):

let's go in little bitty steps

OpenStudy (freckles):

2y=6 so if I multiply 2 on both sides I get 2(2y)=2(6) so 4y=?

OpenStudy (cenaida):

12?

OpenStudy (freckles):

perfect!:) \[3x+4y=9 \\ 3x+12=9 \text{ solve this equation for } x \]

OpenStudy (freckles):

So I replaced 4y with 12

OpenStudy (freckles):

now all have to do is solve 3x+12=9 for x Here is one step 3x=9-12

OpenStudy (cenaida):

what do i do?

OpenStudy (freckles):

you are solving for x

OpenStudy (freckles):

you had 3x+12=9 I subtracted 12 on both sides for you and gave you 3x=9-12

OpenStudy (freckles):

find the difference of 9 and 12 then divide both sides by 3 to get x by itself

OpenStudy (cenaida):

9*3=27 12*3=36

OpenStudy (freckles):

9-12 is -3 isn't it?

OpenStudy (cenaida):

yes

OpenStudy (freckles):

3x+12=9 Subtracted 12 on both sides 3x=9-12 so then 3x=-3

OpenStudy (freckles):

now I actually gave you the last step to perform which was to divide both sides by 3

OpenStudy (cenaida):

9?

OpenStudy (freckles):

are you not wanting to divide?

OpenStudy (freckles):

\[3x=-3 \\ \text{ divide both sides by 3 } \\ \frac{3x}{3}=\frac{-3}{3} \\ x=\frac{-3}{3}\]

OpenStudy (freckles):

how many times does 3 go into -3?

OpenStudy (cenaida):

1

OpenStudy (freckles):

well -1 times

OpenStudy (freckles):

-3/3 is -1

OpenStudy (freckles):

not 9 or -9

OpenStudy (freckles):

-3(3)=-9 but we had -3/3

OpenStudy (cenaida):

ok..thanks so much!